OpenStudy (kirbykirby):

If T~BIN(n, p), then find the expectation: E[T(T-1)...(T-k+1)]

OpenStudy (kirbykirby):

Basically, this is trying to compute the sum \[\sum_{t=0}^n t(t-1)(t-2)\cdots (t-k+1) {n \choose t} p^t (1-p)^{n-t} \] I tried doing this: \[\sum_{t=k}^n {n \choose t} p^t (1-p)^{n-t} \]since the sum is 0 at \(t=0, 1, \ldots, k-1\)\[=\sum_{t=0}^{n-k} {n \choose t+k} p^{t+k} (1-p)^{n-(t+k)} \]\[=p^k\sum_{t=0}^{n-k} {n \choose t+k} p^{t} (1-p)^{(n-k)-t} \] I feel like I need to change \({n \choose t+k}\) to \({n-k \choose t}\) somehow so that I get a Binomial(n-k, p) and then that would sum to 1..but I am having trouble doing so.

OpenStudy (anonymous):

Wait a minute, what is k?

OpenStudy (anonymous):

Just some constant integer you have to deal with?

OpenStudy (anonymous):

What you are getting is something similar to the CDF for the binomial distribution. I don't think there is any simplifying it past that.

OpenStudy (kirbykirby):

Yes k is just some integer with values 1, 2, ...

OpenStudy (anonymous):

On second though, when \(t \geq k\) the \(t(t-1)...(t-k+1)\) will not necessarily be \(1\).

OpenStudy (anonymous):

Recall the derivation the expectation of \(E(T)\): \[\begin{align*}E(T)&=\sum_{t=0}^nt\binom nt p^t (1-p)^{n-t}\\ &=\sum_{t=0}^nt\frac{n!}{t!(n-t)!} p^t (1-p)^{n-t}\\ &=0+\sum_{t=1}^nt\frac{n!}{t!(n-t)!} p^t (1-p)^{n-t}\\ &=np\sum_{t=1}^{n-1}\frac{(n-1)!}{(t-1)!(n-t)!} p^{t-1} (1-p)^{n-t}\\ &=np\sum_{t=1}^{n-1}\frac{(n-1)!}{(t-1)!((n-1)-(t-1))!} p^{t-1} (1-p)^{(n-1)-(t-1)}\\ &=np\sum_{t=0}^{n-1}\frac{(n-1)!}{t!((n-1)-t)!} p^{t-1} (1-p)^{(n-1)-t}\\ &=np\color{red}{\sum_{t=0}^{n-1}\binom{n-1}t p^{t-1} (1-p)^{(n-1)-t}}\\ &=np \end{align*}\] The red part is 1, since that's the pdf of \(\text{Bin}(n-1,p)\). No doubt finding the expectation of \(E\bigg[T(T-1)\cdots(T-(k-1))\bigg]\) involves a similar process.

OpenStudy (anonymous):

When \(t=k\) you have \(t(t−1)...(t−k+1) = k(k-1)\dots (1) = k!\) When \(t=k+1\) you get \(\frac{(k+1)!}{1!}\). Whne \(t=k+c\) you get \((n+c)!/c!\)

OpenStudy (anonymous):

whoops, should say \((k+c)!/c!\)

OpenStudy (anonymous):

And \(c\) is just \(t-k\). So that simplifies to \(t(t−1)...(t−k+1) = t!/(t-k)!\). This is permutations formula, and is just \(k! \binom tk\).

OpenStudy (anonymous):

Hopefully, someone can check if I'm right: \[E\bigg[T(T-1)\cdots(T-(k-1))\bigg]\\ ~~~~=\sum_{t=0}^nt(t-1)\cdots(t-(k-1))\binom nt p^t (1-p)^{n-t}\\ ~~~~=\sum_{t=0}^n t (t-1)\cdots(t-(k-1))\frac{n!}{t!(n-t)!} p^t (1-p)^{n-t}\\ ~~~~=0+\sum_{t=k}^n t (t-1)\cdots(t-(k-1))\frac{n!}{t!(n-t)!} p^t (1-p)^{n-t}\\ ~~~~=n(n-1)\cdots(n-(k-1))~p^{k-2}\\ ~~~~~~~~\times \sum_{t=k}^n t \frac{(n-(k-2))!}{(t-(k-2))!(n-(k-2)-(t-(k-2)))!} p^{t-(k-2)} (1-p)^{n-(k-2)-(t-(k-2))}\\ ~~~~=n(n-1)\cdots(n-(k-1))~p^{k-2}\\ ~~~~=p^{k-2}\frac{n!}{(n-k)!}\] As @wio mentioned, the fraction is equivalent to \({}_nP_k\).